A team played 5 more games this season, than last season. Last season, they won 60% of their games. This season, they won 55% of their games. The team won the same number of games, both seasons. Find the number of games played, in each season.
Answer by oberobic(2304)

x = games played last season
.6*x = games won last season
.
y = games played this season
.55*y = games won this season
y = x+5
.
.6x = .55y
.
.6x = .55(x+5)
.
multiply by 100 to eliminate decimals
.
60x = 55(x+5)
.
60x = 55x + 275
.
5x = 275
.
x = 55 games played last season
.
x+5 = 60 games played this season
.
Always check your answer.
.
.6 * 55 = 33 games won
.55 * 60 = 33 games won
Correct.